The 90% confidence interval for a population mean is (1.2, 5.2).
A. Then the population mean is 3.2, and the margin of error is 2.
B. Then the sample mean is 3.2, and the margin of error is 2.
C. Then the sample mean is 2, and the margin of error is 3.2.
D. Then the population mean is 2, and the margin of error is 3.2.

Answer by stanbon(75887)

The 90% confidence interval for a population mean is (1.2, 5.2).
ME = (1/2)(5.2-1.2) = (1/2)(4) = 2
---
Mean = (5.2+1.2)/3 = 6.4/2 = 3.2
--------------
Answer: A
----
Cheers,
Stan H.
================================
A. Then the population mean is 3.2, and the margin of error is 2.
B. Then the sample mean is 3.2, and the margin of error is 2.
C. Then the sample mean is 2, and the margin of error is 3.2.
D. Then the population mean is 2, and the margin of error is 3.2.